# PASTA: A Scalable Framework for Multi-Policy AI Compliance Evaluation

### v1

@@ 1. Introduction
Current research prototypes for AI policy compliance evaluation are often are tailored to a single major policy  (Dobrescu et al., 2024; Stanford HAI, 2023) or rely on complex technical input requirements  (Kim et al., 2024; Ahmad et al., 2023; Madaio et al., 2024). Simliarly, commercial solutions typically approach AI governance using: (1) policy-specific questionnaires (e.g., LumeNova, VerifyWise, Collibra AI, Credo AI), which hard-code regulations into survey logic and require one survey per policy, offering no flexible or unified input format  (Lum, ; Ver, ; Col, 2024; Credo AI, 2025); and (2) code-scanning or automated workflows (e.g., OneTrust, OpenLayer, Vanta), which depend on access to system metadata or source code and therefore only support compliance once a system is already implemented  (OneTrust, 2025; Openlayer, 2025; Vanta, Inc., 2025). Their high costs also make them inaccessible to individual practitioners and smaller teams. These approaches leave gaps in flexibility, early-stage usability, and scalablity. To address these limitations, we introduce PASTA (Policy Aggregator & Scanner for Trustworthy AI), an AI compliance evaluation tool that leverages LLMs to assess AI system documentation across multiple global policies, including the EU AI Act, AIDA, GDPR, CCPA, and the Colorado AI Act. PASTA is designed for individual practitioners (e.g., developers, project managers, and UI designers) or small teams who typically (1) lack expertise in AI policy compliance and legal knowledge, and (2) operate with limited resources. PASTA provides a multi-policy framework with relatively low-effort input design and interpretable compliance outputs, making AI policy evaluation more accessible and actionable. Specifically, our system addresses the existing gap in the compliance evaluation tool landscape by:
@@
To assess the usability and accuracy of PASTA, we conducted both a two-part user study with 12 AI practitioners and a structured expert-based technical evaluation. Results showed that PASTA enabled comprehensive multi-policy assessments while keeping the entry barrier manageable: participants completed the model card input in an average of 28.4 minutes ($SD=6.7$) and described the process as cognitively effortful yet highly valuable for reflecting on system design and documentation. The resulting compliance reports were viewed as actionable and interpretable, with users confidently identifying policy-specific risks ($M=4.73$, $SD=0.45$) and linking them to system components in under seven minutes on average. Our technical evaluation demonstrated that PASTA’s violation and relevance scores closely aligned with expert consensus (Spearman $\rho=0.6264$ and $\rho=0.7611$), with mean absolute errors remaining low and 87–94% of predictions falling within one point of expert labels. Taken together, these findings indicate that PASTA not only supports scalable and interpretable compliance evaluation but also produces outputs that are directionally consistent with legal expert reasoning, shaping practitioners’ understanding of AI compliance and prompting earlier integration of compliance considerations in their workflows.
@@ 2.2. Existing Landscape of Automated Tools within AI Governance (Algorithmic Auditing)
A growing body of research has emerged to address the challenges of ensuring compliance in AI systems. Building on a recent review synthesizing the breadth of research supporting AI governance  (Batool et al., 2023; Lu et al., 2024), we group existing work into two main approaches: conceptual methodologies and automated tools. Conceptual methodologies include governance frameworks, ethical principles, policies, and high-level guidelines that offer theoretical direction but often lack enforcement mechanisms  (Raji et al., 2020; Kawakami et al., 2024). In contrast, Automated tools instead translate requirements into executable checks and algorithmic auditing. Our work situates within automated AI governance tools, emphasizing on enhancing accessibility and scalability for AI practitioners without legal expertise or support. To structure our review, we further classify existing automated tools into four categories: Risk Prediction Tools, System Behavior Verification Tools, Policy Interpreters, and Commercial AI Governance Platforms.
@@
Efforts to summarize, transform, and make complex legal texts actionable have led to AI governance tools designed to help practitioners more effectively navigate compliance requirements.  (Mittal et al., 2024; Arnold et al., 2024; Constantinides et al., 2024). Many use LLMs and human-centered methods to interpret regulations and generate concise summaries. For example, Mittal et al. automate GDPR privacy-policy interpretation to extract PII-related requirements and produce compliance-oriented summaries  (Mittal et al., 2024). Meanwhile, Arnold et al.’s AGORA dataset offers a structured, cross-jurisdictional taxonomy of AI laws to support regulatory comparison  (Arnold et al., 2024). Although such tools make regulatory information more accessible, they focus primarily on summarization and provide limited compliance evaluation features.
@@
A broad ecosystem of commercial AI governance platforms has emerged to support organizations in managing AI systems across their lifecycle. These platforms are designed primarily for enterprise regulatory workflows, where teams coordinate model registration, risk assessment, testing, documentation  (Credo AI, 2025; Col, 2024; OneTrust, 2025; Ver, ; Lum, ; Vanta, Inc., 2025; Openlayer, 2025; Amazon Web Services, 2025; Atlan Technologies, Inc., 2024; IBM, 2025). Our analysis of 10 prominent platforms shows that, despite differences in emphasis, they converge around a common set of functional components:
@@
Although existing platforms offer mature all-in-one governance infrastructures for enterprises, they typically keep their implementations confidential due to commercial value, limiting their contribution to research or community advancement in AI governance. Their pricing models also reflect their enterprise orientation, with many solutions requiring custom sales engagements and annual costs that place them out of reach for individual practitioners, researchers, and small organizations who typically (1) lack specialized expertise in AI policy compliance and legal requirements, and (2) operate with limited resources, making affordable automated tools a essential alternative rather than costly human evaluators (see Table  1 for pricing).
For the subset of platforms that include regulatory compliance assessments, the component most relevant to our work, we identify 2 dominant approaches and the gaps they introduce:
@@
  • Code-Scanning and Automated Compliance Approaches (OneTrust, OpenLayer, Vanta). These platforms embed compliance checks into automated technical workflows, such as CI/CD pipelines or continuous monitoring systems. For example, OneTrust scans websites and applications for privacy and regulatory risks, maps them to relevant policies, and provides remediation guidance. As these automated solutions require access to system metadata or srouce code, they can only be used upon implemented system. This leaves a gap for flexible solutions that support compliance assessment earlier in the development process.
@@ 2.3. AI Documentation Standards and Model Cards
The EU AI Act-Oriented Extensions of Model Cards
Several recent extensions explicitly center on alignment with the EU AI Act. For example, Marino et al. introduced compliance Cards as a machine-readable format that encodes EU AI Act-specific metadata as an input format to a rules-based Algorithm to generate compliance predictions  (Marino et al., 2024). Similarly, Use Case Cards standardizes descriptions of an AI system by combining UML diagrams with structured tables to articulate real-world scenarios and anticipate obligations under the Act’s risk-based framework  (Bogucka et al., 2024). These artifacts implement EU AI Act requirements and are designed around that specific regulatory context.
@@
Some documenting approaches focus on technical governance. Policy Cards introduce a machine-readable, deployment-layer artifact that specifies what autonomous AI agents are permitted, required, or forbidden to do at runtime. Encoded as JSON schemas, Policy Cards can be integrated with CI/CD pipelines and map operational constraints to major assurance frameworks such as the NIST AI RMF, ISO/IEC 42001, and the EU AI Act, focusing on regulating system behavior at runtime and addressing the technical enforcement layer.  (Mavracic, 2025).
@@ 3. Design Goals
Prior work has advanced automated AI governance support through various approaches. Yet our review identifies several persistent gaps. First, most questionnaire-based tools remain hard-coded a single regulatory regime  (Ver, ; Lum, ; Credo AI, 2025; Col, 2024), limiting their flexibility in today’s rapidly expanding policy landscape (OECD, 2023; Stanford Institute for Human-Centered Artificial Intelligence, 2024). Second, code-scanning solutions are only available to implemented systems, leaving early- and mid-stage projects without support  (Openlayer, 2025; OneTrust, 2025; Vanta, Inc., 2025; Ahmad et al., 2023). Finally, commercial AI governance platforms introduce substantial financial barriers and keep their implementation details confidential (Table 1), placing these workflows out of reach for individual practitioners and small teams.
Guided by these gaps and informed by research on AI documentation (Mitchell et al., 2019; Bogucka et al., 2025; Marino et al., 2024) and LLM-based legal reasoning (Wang et al., 2025; Garza et al., 2024), we derive 3 design goals for PASTA:
  • Design Goal 1: Scalability and Adaptability Across Multiple Diverse Policies. PASTA should evalute heterogeneous, cross-jurisdictional policies, motivated by the rapid growth of global AI regulation  (OECD, 2023; Stanford Institute for Human-Centered Artificial Intelligence, 2024), and the limitation of existing single-policy tools  (Ver, ; Credo AI, 2025).
@@ 4. PASTA: Policy Aggregator & Scanner for Trustworthy AI
To address our research questions, we developed PASTA and conducted assessments to evaluate its accuracy, scalability and usability. This section details our solution.
@@ 4.2. System Architecture Overview
PASTA is built on a modular architecture that integrates several interconnected components to facilitate AI policy compliance evaluation (Fig. 1 depicts the PASTA workflow). The system features:
@@
The overall evaluation pipeline in PASTA follows a model card-to-policy comparison workflow. This process begins by taking the user-uploaded model card, structured by sections, and comparing it against preprocessed policy documents such as AIDA, CCPA, and the EU AI Act, which are segmented in articles. To optimize processing, an irrelevancy filter first eliminates unrelated pairs of model card sections and policy articles to reduce computational load. The remaining relevant pairs are then evaluated using a large language model (LLM) that performs pairwise comparisons to assess potential policy violations and generate structured reasoning. These results are compiled into a dataset, which was subsequently aggregated into human-readable summaries at multiple granularity levels. Finally, the summarized results are displayed on the user interface through interactive heatmaps, tables, and text blocks to support interpretable, actionable compliance assessments for AI practitioners.
@@ 4.3. Designing a Model Card Template Specialized for Multi-Policy Compliance
  • Example: The General Information category contains section System Name, Versioning Details, Primary Developer, Contact Information, and System Overview.
    See Table. 2 for the complete model card structure.
@@
  • Training Data Information: Combines the Evaluation Data and Training Data sections from Mitchell et al. and the Kaggle Model Card  (Mitchell et al., 2019; Shankar, 2023). This unified structure is better suited for AI systems beyond machine learning models, where the distinction between training and evaluation datasets is often less meaningful. By consolidating these sections, we emphasize a system-level perspective on data provenance, collection practices, and bias mitigation, aligning more closely with the needs of compliance evaluation and regulatory transparency.
@@ 4.4. Policy Normalization and Preprocessing for Scalable Evaluation
By treating the paragraph as the atomic unit of analysis across all jurisdictions, PASTA processes any policy through a single pipeline, allowing new regulations to be added without changing the evaluation logic. We standardize policy content into a Markdown table that labels each paragraph by policy name, article number, and paragraph index, creating a uniform structure aligned with the PASTA model card and supporting scalable policy–system comparison (see Fig. 2 and Table. 3). An LLM-powered preprocessing step automatically extracts and formats paragraph-level units from raw HTML, enabling consistent structuring of long, heterogeneous policy documents and ensuring reliable downstream alignment with model card content.
@@ 4.5. The Pairwise Compliance Evaluation Engine
For each comparison pair, an LLM assesses whether the model card section satisfies, partially addresses, or violates the policy article’s requirements. The full model card and policy context are cached and provided with each request. Figure 3 illustrates how comparison pairs are structured and batched within prompts. Pairs previously flagged as irrelevant (Section 4.4) are skipped to reduce unnecessary computation.
We have selected Claude Sonnet 4 as the evaluation model, and the selection process is detailed in the Appendix A.1.
@@
Each comparison pair receives a violation score from 0 to 5, where 0 indicates full compliance and 5 signals a critical violation. This scale is expressive enough to distinguish minor ambiguities from serious risks while remaining intuitive for practitioners without legal training (Alok Joshi, 2021). See Table. 13 in Appendix for the violation score rubric.

### v2

@@ Abstract.
AI governance, AI compliance, policy evaluation, model cards, large language models, human-centered AI
copyright: acmlicensedjournalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790966isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Interactive systems and toolsccs: Computing methodologies Artificial intelligenceccs: Social and professional topics Computing / technology policy
@@ 1. Introduction
Current research prototypes for AI policy compliance evaluation are often tailored to a single major policy  (Dobrescu et al., 2024; Stanford HAI, 2023) or rely on complex technical input requirements  (Kim et al., 2024; Ahmad et al., 2023; Madaio et al., 2024). Similarly, commercial solutions typically approach AI governance using: (1) policy-specific questionnaires (e.g., LumeNova, VerifyWise, Collibra AI, Credo AI), which hard-code regulations into survey logic and require one survey per policy, offering no flexible or unified input format  (Lum, [n. d.]; Ver, [n. d.]; Col, 2024; Credo AI, 2025); and (2) code-scanning or automated workflows (e.g., OneTrust, OpenLayer, Vanta), which depend on access to system metadata or source code and therefore only support compliance once a system is already implemented  (OneTrust, 2025; Openlayer, 2025; Vanta, Inc., 2025). Their high costs also make them inaccessible to individual practitioners and smaller teams. These approaches leave gaps in flexibility, early-stage usability, and scalability. To address these limitations, we introduce PASTA (Policy Aggregator & Scanner for Trustworthy AI), an AI compliance evaluation tool that leverages LLMs to assess AI system documentation across multiple global policies, including the EU AI Act, AIDA, GDPR, CCPA, and the Colorado AI Act. PASTA is designed for individual practitioners (e.g., developers, project managers, and UI designers) or small teams who typically (1) lack expertise in AI policy compliance and legal knowledge, and (2) operate with limited resources. PASTA provides a multi-policy framework with relatively low-effort input design and interpretable compliance outputs, making AI policy evaluation more accessible and actionable. Specifically, our system addresses the existing gap in the compliance evaluation tool landscape by:
@@
To assess the usability and accuracy of PASTA, we conducted both a two-part user study with 12 AI practitioners and a structured expert-based technical evaluation. Results showed that PASTA enabled comprehensive multi–policy assessments while keeping the entry barrier manageable: participants completed the model card input in an average of 28.4 minutes ($\mathrm{SD}=6.7$) and described the process as cognitively effortful yet highly valuable for reflecting on system design and documentation. The resulting compliance reports were viewed as actionable and interpretable, with users confidently identifying policy-specific risks ($\mathrm{M}=4.73$, $\mathrm{SD}=0.45$) and linking them to system components in under seven minutes on average. Our technical evaluation demonstrated that PASTA’s violation and relevance scores closely aligned with expert consensus (Spearman $\rho=0.6264$ and $\rho=0.7611$), with mean absolute errors remaining low and 87–94% of predictions falling within one point of expert labels. Taken together, these findings indicate that PASTA not only supports scalable and interpretable compliance evaluation but also produces outputs that are directionally consistent with legal expert reasoning, shaping practitioners’ understanding of AI compliance and prompting earlier integration of compliance considerations in their workflows.
@@ 2.2. Existing Landscape of Automated Tools within AI Governance (Algorithmic Auditing)
A growing body of research has emerged to address the challenges of ensuring compliance in AI systems. Building on a recent review synthesizing the breadth of research supporting AI governance  (Batool et al., 2023; Lu et al., 2024), we group existing work into two main approaches: conceptual methodologies and automated tools. Conceptual methodologies include governance frameworks, ethical principles, policies, and high-level guidelines that offer theoretical direction but often lack enforcement mechanisms  (Raji et al., 2020; Kawakami et al., 2024). In contrast, automated tools instead translate requirements into executable checks and algorithmic auditing. Our work situates within automated AI governance tools, emphasizing enhancing accessibility and scalability for AI practitioners without legal expertise or support. To structure our review, we further classify existing automated tools into four categories: Risk Prediction Tools, System Behavior Verification Tools, Policy Interpreters, and Commercial AI Governance Platforms.
@@
Efforts to summarize, transform, and make complex legal texts actionable have led to AI governance tools designed to help practitioners more effectively navigate compliance requirements  (Mittal et al., 2024; Arnold et al., 2024; Constantinides et al., 2024). Many use LLMs and human-centered methods to interpret regulations and generate concise summaries. For example, Mittal et al. automate GDPR privacy-policy interpretation to extract PII-related requirements and produce compliance-oriented summaries  (Mittal et al., 2024). Meanwhile, Arnold et al.’s AGORA dataset offers a structured, cross-jurisdictional taxonomy of AI laws to support regulatory comparison  (Arnold et al., 2024). Although such tools make regulatory information more accessible, they focus primarily on summarization and provide limited compliance evaluation features.
@@
A broad ecosystem of commercial AI governance platforms has emerged to support organizations in managing AI systems across their lifecycle. These platforms are designed primarily for enterprise regulatory workflows, where teams coordinate model registration, risk assessment, testing and documentation  (Credo AI, 2025; Col, 2024; OneTrust, 2025; Ver, [n. d.]; Lum, [n. d.]; Vanta, Inc., 2025; Openlayer, 2025; Amazon Web Services, 2025; Atlan Technologies, Inc., 2024; IBM, 2025). Our analysis of 10 prominent platforms shows that, despite differences in emphasis, they converge around a common set of functional components:
@@
Although existing platforms offer mature all-in-one governance infrastructures for enterprises, they typically keep their implementations confidential due to commercial value, limiting their contribution to research or community advancement in AI governance. Their pricing models also reflect their enterprise orientation, with many solutions requiring custom sales engagements and annual costs that place them out of reach for individual practitioners, researchers, and small organizations who typically (1) lack specialized expertise in AI policy compliance and legal requirements, and (2) operate with limited resources, making affordable automated tools an essential alternative rather than costly human evaluators (see Table  1 for pricing).
For the subset of platforms that include regulatory compliance assessments, the component most relevant to our work, we identify two dominant approaches and the gaps they introduce:
@@
  • Code-Scanning and Automated Compliance Approaches (OneTrust, OpenLayer, Vanta). These platforms embed compliance checks into automated technical workflows, such as CI/CD pipelines or continuous monitoring systems. For example, OneTrust scans websites and applications for privacy and regulatory risks, maps them to relevant policies, and provides remediation guidance. As these automated solutions require access to system metadata or source code, they can only be used on implemented systems. This leaves a gap for flexible solutions that support compliance assessment earlier in the development process.
@@ 2.3. AI Documentation Standards and Model Cards
The EU AI Act–Oriented Extensions of Model Cards
Several recent extensions explicitly center on alignment with the EU AI Act. For example, Marino et al. introduced compliance Cards as a machine-readable format that encodes EU AI Act-specific metadata as an input format to a rules-based algorithm to generate compliance predictions  (Marino et al., 2024). Similarly, Use Case Cards standardize descriptions of an AI system by combining UML diagrams with structured tables to articulate real-world scenarios and anticipate obligations under the Act’s risk-based framework  (Bogucka et al., 2024). These artifacts implement EU AI Act requirements and are designed around that specific regulatory context.
@@
Some documenting approaches focus on technical governance. Policy Cards introduce a machine-readable, deployment-layer artifact that specifies what autonomous AI agents are permitted, required, or forbidden to do at runtime. Encoded as JSON schemas, Policy Cards can be integrated with CI/CD pipelines and map operational constraints to major assurance frameworks such as the NIST AI RMF, ISO/IEC 42001, and the EU AI Act, focusing on regulating system behavior at runtime and addressing the technical enforcement layer  (Mavracic, 2025).
@@ 3. Design Goals
Prior work has advanced automated AI governance support through various approaches. Yet our review identifies several persistent gaps. First, most questionnaire-based tools remain hard-coded to a single regulatory regime  (Ver, [n. d.]; Lum, [n. d.]; Credo AI, 2025; Col, 2024), limiting their flexibility in today’s rapidly expanding policy landscape (OECD, 2023; Stanford Institute for Human-Centered Artificial Intelligence, 2024). Second, code-scanning solutions are only available to implemented systems, leaving early- and mid-stage projects without support  (Openlayer, 2025; OneTrust, 2025; Vanta, Inc., 2025; Ahmad et al., 2023). Finally, commercial AI governance platforms introduce substantial financial barriers and keep their implementation details confidential (Table  1), placing these workflows out of reach for individual practitioners and small teams.
Guided by these gaps and informed by research on AI documentation (Mitchell et al., 2019; Bogucka et al., 2025; Marino et al., 2024) and LLM-based legal reasoning (Wang et al., 2025; Garza et al., 2024), we derive three design goals for PASTA:
  • Design Goal 1: Scalability and Adaptability Across Multiple Diverse Policies. PASTA should evaluate heterogeneous, cross-jurisdictional policies, motivated by the rapid growth of global AI regulation  (OECD, 2023; Stanford Institute for Human-Centered Artificial Intelligence, 2024), and the limitation of existing single-policy tools  (Ver, [n. d.]; Credo AI, 2025).
@@ 4. PASTA: Policy Aggregator & Scanner for Trustworthy AI
To address our research questions, we developed PASTA and conducted assessments to evaluate its accuracy, scalability, and usability. This section details our solution.
@@ 4.2. System Architecture Overview
PASTA is built on a modular architecture that integrates several interconnected components to facilitate AI policy compliance evaluation (Fig.  1 depicts the PASTA workflow). The system features:
@@
The overall evaluation pipeline in PASTA follows a model card-to-policy comparison workflow. This process begins by taking the user-uploaded model card, structured by sections, and comparing it against preprocessed policy documents such as AIDA, CCPA, and the EU AI Act, which are segmented into articles. To optimize processing, an irrelevancy filter first eliminates unrelated pairs of model card sections and policy articles to reduce computational load. The remaining relevant pairs are then evaluated using a large language model (LLM) that performs pairwise comparisons to assess potential policy violations and generate structured reasoning. These results are compiled into a dataset, which was subsequently aggregated into human-readable summaries at multiple granularity levels. Finally, the summarized results are displayed on the user interface through interactive heatmaps, tables, and text blocks to support interpretable, actionable compliance assessments for AI practitioners.
@@ 4.3. Designing a Model Card Template Specialized for Multi-Policy Compliance
  • Example: The General Information category contains sections System Name, Versioning Details, Primary Developer, Contact Information, and System Overview.
    See Table.  2 for the complete model card structure.
@@
  • System Data Information: Combines the Evaluation Data and Training Data sections from Mitchell et al. and the Kaggle Model Card  (Mitchell et al., 2019; Shankar, 2023). This unified structure is better suited for AI systems beyond machine learning models, where the distinction between training and evaluation datasets is often less meaningful. By consolidating these sections, we emphasize a system-level perspective on data provenance, collection practices, and bias mitigation, aligning more closely with the needs of compliance evaluation and regulatory transparency.
@@ 4.4. Policy Normalization and Preprocessing for Scalable Evaluation
By treating the paragraph as the atomic unit of analysis across all jurisdictions, PASTA processes any policy through a single pipeline, allowing new regulations to be added without changing the evaluation logic. We standardize policy content into a Markdown table that labels each paragraph by policy name, article number, and paragraph index, creating a uniform structure aligned with the PASTA model card and supporting scalable policy–system comparison (see Fig.  2 and Table. 3). An LLM-powered preprocessing step automatically extracts and formats paragraph-level units from raw HTML, enabling consistent structuring of long, heterogeneous policy documents and ensuring reliable downstream alignment with model card content.
@@ 4.5. The Pairwise Compliance Evaluation Engine
For each comparison pair, an LLM assesses whether the model card section satisfies, partially addresses, or violates the policy article’s requirements. The full model card and policy context are cached and provided with each request. Fig.  3 illustrates how comparison pairs are structured and batched within prompts. Pairs previously flagged as irrelevant (Section  4.4) are skipped to reduce unnecessary computation.
We have selected Claude Sonnet 4 as the evaluation model, and the selection process is detailed in the Appendix  A.1.
@@
Each comparison pair receives a violation score from 0 to 5, where 0 indicates full compliance and 5 signals a critical violation. This scale is expressive enough to distinguish minor ambiguities from serious risks while remaining intuitive for practitioners without legal training  (Alok Joshi, 2021). See Table.  12 in Appendix for the violation score rubric.